OpenStudy (anonymous):

X^2-11x+28=0 Don't know where to start

OpenStudy (anonymous):

first factor out (x-7)(x-4)=0 x=7 and x=4

OpenStudy (anonymous):

then would you just make x-7=0; X-4=0

OpenStudy (anonymous):